# definition

### v2

@@ Abstract
Heterogeneity poses a fundamental challenge for many real-world large-scale decision-making problems but remains largely understudied. In this paper, we study the fully heterogeneous setting of a prominent class of such problems, known as weakly-coupled Markov decision processes (WCMDPs). Each WCMDP consists of $N$ arms (or subproblems), which have distinct model parameters in the fully heterogeneous setting, leading to the curse of dimensionality when $N$ is large. We show that, under mild assumptions, a natural adaptation of the ID policy, although originally proposed for a homogeneous special case of WCMDPs, in fact achieves an $O(1/\sqrt{N})$ optimality gap in long-run average reward per arm for fully heterogeneous WCMDPs as $N$ becomes large. This is the first asymptotic optimality result for fully heterogeneous average-reward WCMDPs. Our techniques highlight the construction of a novel projection-based Lyapunov function, which witnesses the convergence of rewards and costs to an optimal region in the presence of heterogeneity.
Keywords: weakly-coupled Markov decision processes, fully heterogeneous systems, asymptotic optimality, planning, average-reward Markov decision processes, Lyapunov analysis
@@ 1 Introduction
Heterogeneity poses a fundamental challenge for many real-world decision-making problems, where each problem consists of a large number of interacting components. However, despite its practical significance, heterogeneity remains largely understudied in the literature. In this paper, we study heterogeneous settings of a prominent class of such problems, known as weakly-coupled Markov decision processes (WCMDPs) [Hawkins, 2003]. A WCMDP consists of $N$ arms (or subproblems), where each arm itself is a Markov decision process (MDP). In a heterogeneous setting, the arms have distinct model parameters. At each time step, the decision-maker selects an action for each arm, which affects the arm’s transition probabilities and reward, and then the arms make state transitions independently. However, these actions are subject to a set of global budget constraints, where each constraint limits one type of total cost across all arms at each time step. The objective is to find a policy that maximizes the long-run average reward over an infinite time horizon. We focus on the planning setting, where all the model parameters are known.
WCMDPs have been used to model a wide range of applications, including online advertising [Boutilier and Lu, 2016, Zhou et al., 2023], job scheduling [Yu et al., 2018], healthcare [Biswas et al., 2021], surveillance [Villar, 2016], and machine maintenance [Glazebrook et al., 2005]. A faithful modeling of these applications calls for heterogeneity. For instance, in [Biswas et al., 2021], arms are beneficiaries of a health program and they could react to interventions differently; in [Villar, 2016], arms are targets of surveillance who have different locations and probabilities to be exposed; in [Glazebrook et al., 2005], arms are machines that could require distinct repair schedules.
Although heterogeneity is crucial in the modeling of these applications, most existing work on average-reward WCMDPs (or their special cases) establish asymptotic optimality only for the homogeneous setting where all arms share the same set of model parameters [Weber and Weiss, 1990, Verloop, 2016, Gast et al., 2023a, b, Hong et al., 2023, 2024a, 2024b, Yan, 2024, Goldsztajn and Avrachenkov, 2024]. An exception is [Verloop, 2016], which considers the typed heterogeneous setting, where the $N$ arms are divided into a constant number of types as $N$ scales up, with each type having distinct model parameters. While heterogeneous WCMDPs have been studied under the finite-horizon total-reward and discounted-reward criteria, as we review in Appendix A, these results do not extend to the average-reward setting we consider.
@@
Our contribution
In this paper, we study fully heterogeneous WCMDPs. We consider a natural adaptation of the ID policy, originally proposed by Hong et al. [2024a] for homogeneous restless bandits—a renowned special case of WCMDPs. Given that the ID policy was designed for a much simpler setting, its effectiveness in fully heterogeneous WCMDPs is far from obvious. However, with proper adaptation and novel theoretical techniques, we show that the ID policy achieves an $O(1/\sqrt{N})$ optimality gap under mild assumptions as the number of arms $N$ becomes large. Here, the optimality gap is the gap between the long-run average reward per arm achieved by the ID policy and that under the optimal policy. This is the first result establishing asymptotic optimality for fully heterogeneous average-reward WCMDPs.
@@
Here we briefly describe how the adapted ID policy works. While it retains the core structure of the original ID policy, modifications are needed to handle heterogeneity. The policy consists of two phases. The first phase is a pre-processing phase, where we compute an optimal single-armed policy for each arm (denoted as ${{\bar{\pi}}^{*}}_{i}$ for the $i$-th arm) that prescribes the ideal action the arm would take at each state. These optimal single-armed policies can be efficiently solved via a linear program. The second phase is the real-time phase. At each time step, the policy iterates over the arms in a fixed order, allowing as many arms as possible to follow their respective ideal actions while satisfying the budget constraints. Unlike the original ID policy, which has only one optimal single-armed policy due to homogeneous arms, our adapted version computes $N$ optimal single-armed policies, one for each arm. Additionally, we introduce an ID reassignment procedure before the real-time phase that reorders arms to ensure a regularity property.
Our assumptions are in terms of the optimal single-armed policies. We assume that they induce aperiodic unichains and their transition probability matrices have spectral gaps bounded away from zero. These assumptions generalize the one in [Hong et al., 2023] and are weaker than most assumptions in previous papers when specialized to their settings.
Technical novelty
@@
The main technical innovation of the paper is the introduction of a novel Lyapunov function for fully heterogeneous WCMDPs. Specifically, to prove the asymptotic optimality of a policy, a key step is to show that the system state is globally attracted to an optimal region where most arms can follow the ideal actions generated by their respective optimal single-armed policies ${{\bar{\pi}}^{*}}_{i}$’s. Let ${\bm{S}}_{t}=(S_{1,t},\dots,S_{N,t})$ denote the joint state of the $N$ arms at time $t$. Then this optimal region consists of those states ${\bm{S}}_{t}$’s whose each coordinate $S_{i,t}$ is approximately an independent sample from a certain optimal state distribution $\mu^{*}_{i}$ for the $i$-th arm, for $i=1,2,\dots,N$. In the homogeneous setting, there is only one optimal state distribution $\mu^{*}$. Consequently, the optimal region in that case is the set of system states whose empirical distribution across coordinates remains sufficiently close to $\mu^{*}$; global attraction to this region could be established by a Lyapunov function that depends on the empirical distribution of ${\bm{S}}_{t}$. In the heterogeneous setting, however, it has been unclear how such a Lyapunov function could be constructed. Intuitively, this Lyapunov function should focus on the collective properties of ${\bm{S}}_{t}$ rather than the states of individual arms, so that the statistical patterns across the coordinates of ${\bm{S}}_{t}$ can be captured. Our technique is to project ${\bm{S}}_{t}$ onto a set of carefully selected feature vectors, and define the Lyapunov function based on these projections. These feature vectors encode the minimal amount of information needed to evaluate the relevant functions of the system state (e.g., instantaneous reward or cost) and predict their future expectations. This projection-based Lyapunov function provides a principled way to measure deviations of the system state from the optimal region in a fully heterogeneous setting. A more detailed discussion of this approach can be found in Section 4.
@@
Beyond WCMDPs, our techniques have the potential to be applied to more general heterogeneous large stochastic systems. Heterogeneity has been a topic of strong interest in these systems, but it is known to be a challenging problem with limited theoretical results. Only recently have there been notable breakthroughs. [Allmeier and Gast, 2022, 2024] extended the popular mean-field analysis to a class of heterogeneous large stochastic systems for the first time, but the results are only for transient distributions. Another line of work [Zhao et al., 2024, Zhao and Mukherjee, 2024] studied heterogeneous load-balancing systems. They first analyzed the transient distributions and then used interchange-of-limits arguments to extend the results to steady state. Our method provides a more direct framework for steady-state analysis and has the potential to generalize to a broader range of heterogeneous stochastic systems.
@@
We first focus on the average-reward criterion. As mentioned earlier, most existing work considers the homogeneous setting. Early work on WCMDPs primarily focuses on a special case known as the restless bandit (RB) problem, where each arm’s MDP has a binary action space (active and passive actions) and there is only one budget constraint that limits the total number of active actions across all arms at each time step. The seminal work by Whittle [1988] introduced the RB problem and the celebrated Whittle index policy, which was later shown to achieve an $o(1)$ optimality gap as $N\to\infty$ under a set of conditions [Weber and Weiss, 1990]. Subsequent work on RBs has focused on designing policies that achieve asymptotic optimality under more relaxed conditions [Verloop, 2016, Hong et al., 2023, 2024a, Yan, 2024], as well as improving the optimality gap to $O(1/\sqrt{N})$ [Hong et al., 2023, 2024a] or $O(\exp(-cN))$ [Gast et al., 2023a, b, Hong et al., 2024b]. Among these papers, only [Verloop, 2016] addresses heterogeneity, but in the typed heterogeneous setting, where the $N$ arms are divided into a constant number of types as $N\to\infty$.
Beyond RBs, work on general average-reward WCMDPs is scarce. The closest to ours is [Verloop, 2016], which considered typed heterogeneous WCMDPs with a single budget constraint and established an $o(1)$ optimality gap. More recently, [Goldsztajn and Avrachenkov, 2024] proved the first asymptotic optimality result for homogeneous WCMDPs, also achieving an $o(1)$ optimality gap.
Under the finite-horizon total-reward or discounted-reward criteria, there has been more work on heterogeneous settings, including both the typed heterogeneous setting [D’Aeth et al., 2023, Ghosh et al., 2023] and, more recently, the fully heterogeneous setting [Brown and Smith, 2020, Brown and Zhang, 2022, 2023, Zhang, 2024]. However, the optimality gap in these papers generally grow super-linearly with the (effective) time horizon, except under restrictive conditions. As a result, it would be difficult to extend these results to the average-reward setting and still achieve asymptotic optimality.
@@
Let $\mathbb{R}$, $\mathbb{N}$, and $\mathbb{N}_{+}$ denote the sets of real numbers, nonnegative integers, and positive integers, respectively. Let $[N]\triangleq\{1,2,\dots,N\}$ for any $N\in\mathbb{N}_{+}$ and $[n_{1}:n_{2}]\triangleq\{n_{1},n_{1}+1,\dots,n_{2}\}$ for $n_{1},n_{2}\in\mathbb{N}_{+}$ with $n_{1}\leq n_{2}$. Let $[0,1]_{N}=\{i/N\colon i\in\mathbb{N},0\leq i/N\leq 1\}$, the set of integer multiples of $1/N$ in $[0,1]$. For a matrix ${A}\in\mathbb{R}^{d\times d}$, we denote its operator norm as $\left\lVert{A}\right\rVert_{p}=\sup_{x\neq\mathbf{0}}\frac{\left\lVert{A}x\right\rVert_{p}}{\left\lVert x\right\rVert_{p}}$, and refer to the $2$-norm as $\left\lVert A\right\rVert$ for simplicity. We use boldface letters to denote matrices, and regular letters to denote vectors and scalars. We write $\mathbb{R}^{\mathbb{S}}$ for the set of real-valued vectors indexed by elements of $\mathbb{S}$, or equivalently, the set of real-valued functions on $\mathbb{S}$; for each $v\in\mathbb{R}^{\mathbb{S}}$, let $v(s)$ to denote its element corresponding to $s\in\mathbb{S}$.
@@ 2 Problem setup
A policy $\pi$ for the $N$-armed problem specifies the action for each of the $N$ arms, in a possibly history-dependent way. Under policy $\pi$, let $S_{i,t}^{\pi}$ denote the state of the $i$th arm at time $t$, and we refer to $\bm{S}^{\pi}_{t}\triangleq(S_{i,t}^{\pi})_{i\in[N]}$ as the system state. Similarly, let $A_{i,t}^{\pi}$ denote the action applied to arm $i$ at time $t$, and we refer to $\bm{A}^{\pi}_{t}\triangleq(A_{i,t}^{\pi})_{i\in[N]}$ as the system action. In this paper, we also use an alternative representation of the system state, denoted as ${\bm{X}}_{t}^{\pi}$ and defined as follows. Let $X_{i,t}^{\pi}=(X_{i,t}^{\pi}(s))_{s\in\mathbb{S}}\in\mathbb{R}^{|\mathbb{S}|}$ be a row vector where the entry corresponding to state $s$ is given by $X_{i,t}^{\pi}(s)=\mathbbm{1}\{S_{i,t}^{\pi}=s\}$; i.e., $X_{i,t}^{\pi}$ is a one-hot row vector whose $s$’s entry is $1$ if $S_{i,t}^{\pi}=s$ and is 00 otherwise. Then let ${\bm{X}}_{t}^{\pi}$ be an $N\times|\mathbb{S}|$ matrix whose $i$th row is $X_{i,t}^{\pi}$. It is easy to see that ${\bm{X}}_{t}^{\pi}$ contains the same information as $\bm{S}^{\pi}_{t}$, and we refer to both of them as the system state. In this paper, we often encounter vectors like $X_{i,t}^{\pi}=(X_{i,t}^{\pi}(s))_{s\in\mathbb{S}}$, whose entries correspond to different states in $\mathbb{S}$. For such vectors, say $u$ and $v$, we use the inner product to write a sum for convenience $\left\langle u,v\right\rangle\triangleq\sum_{s\in\mathbb{S}}u(s)v(s)$. We sometimes omit the superscript $\pi$ when it is clear which policy is being used.
@@
Our objective is to maximize the long-run time-average reward subject to the cost constraints. To be more precise, we follow the treatment for maximizing average reward in [Puterman, 2005]. For any policy $\pi$ and an initial state $\bm{S}_{0}$ of the $N$-armed system, consider the limsup average reward $R^{+}(\pi,\bm{S}_{0})$ and the liminf average $R^{-}(\pi,\bm{S}_{0})$, defined as $R^{+}(\pi,\bm{S}_{0})=\lim\sup_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{N}\sum_{i\in[N]}\mathbb{E}\left[r_{i}(S_{i,t}^{\pi},A_{i,t}^{\pi})\right]$ and $R^{-}(\pi,\bm{S}_{0})=\lim\inf_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{N}\sum_{i\in[N]}\mathbb{E}\left[r_{i}(S_{i,t}^{\pi},A_{i,t}^{\pi})\right]$. If $R^{+}(\pi,\bm{S}_{0})=R^{-}(\pi,\bm{S}_{0})$, then the average reward of policy $\pi$ under initial condition $\bm{S}_{0}$ exists and is defined as
@@
Let the optimal value of this problem be denoted as $R^{*}(N,\bm{S}_{0})$. Note that since the WCMDP is an MDP with finite state and action space, if we replace the $R^{-}(\pi,\bm{S}_{0})$ in the objective (3a) with $R^{+}(\pi,\bm{S}_{0})$, the optimal value stays the same.
@@
Recall that exactly solving the WCMDP problem is PSPACE-hard. In this paper, our goal is to design a policy $\pi$ that is efficiently computable and asymptotically optimal as $N\to\infty$, with the following notion for asymptotic optimality. For any policy $\pi$, we define its optimality gap as $R^{*}(N,\bm{S}_{0})-R^{-}(\pi,\bm{S}_{0})$. We say the policy $\pi$ is asymptotically optimal if as $N\to\infty$,
@@
When we take this asymptotic regime as $N\to\infty$, we keep the number of constraints, $K$, as well as the budget coefficients, $\alpha_{1},\alpha_{2},\dots,\alpha_{K}$, fixed. We assume that the reward functions and cost functions are uniformly bounded, i.e., $\sup_{i\in\mathbb{N}_{+}}\max_{s\in\mathbb{S},a\in\mathbb{A}}|r_{i}(s,a)|\triangleq r_{\max}<\infty$ and $\sup_{i\in\mathbb{N}_{+}}\max_{k\in[K],s\in\mathbb{S},a\in\mathbb{A}}c_{k,i}(s,a)\triangleq c_{\max}<\infty$. This notion for asymptotic optimality is consistent with that in the existing literature (e.g., [Verloop, 2016, Definition 4.11]). We are interested in not only achieving asymptotic optimality but also characterizing the order of the optimality gap.
@@
$\sum_{s^{\prime}\in\mathbb{S},a^{\prime}\in\mathbb{A}}\mathbb{P}_{i}(s\mid s^{\prime},a^{\prime})y_{i}(s^{\prime},a^{\prime})=\sum_{a\in\mathbb{A}}y_{i}(s,a),\quad\forall s\in\mathbb{S},\forall i\in[N],$ (5c)
$\sum_{s^{\prime}\in\mathbb{S},a^{\prime}\in\mathbb{A}}y_{i}(s^{\prime},a^{\prime})=1,\qquad y_{i}(s,a)\geq 0,\quad\forall s\in\mathbb{S},\forall a\in\mathbb{A},\forall i\in[N].$ (5d)
@@
Therefore, to derive an upper bound for the optimality gap, it suffices to upper bound $R_{N}^{\textup{rel}}-R^{-}(\pi,\bm{S}_{0})$, which is the route taken by this paper.
@@
Our LP (5) serves a similar role to the LP used in previous work on restless bandits and WCMDPs with homogeneous arms [see, e.g., Weber and Weiss, 1990, Gast et al., 2023a, Hong et al., 2023]. Both our LP and the LP in previous work relax the hard budget constraints to time-average constraints. However, in the homogeneous arm setting, the LP has only one set of state-action frequencies $y(s,a)$, and the LP is independent of $N$. As a result, both the optimal value of the LP and the complexity of solving it are independent of $N$. Some existing work [Verloop, 2016] considers heterogeneous arms, but only in the limited sense of having a constant number of arm types. This setting still closely resembles the homogeneous setting, and the LP remains independent of $N$.
@@
3 ID policy (with reassignment)
In this section, we introduce our adapted version of the ID policy that was originally proposed in [Hong et al., 2024a]. This adapted version retains the core structure of the original ID policy but includes an additional ID reassignment procedure. For simplicity, we will continue to refer to this adapted version as the ID policy. We begin by introducing a building block of the ID policy, referred to as optimal single-armed policies, followed by a brief description of the ID reassignment algorithm. We then present the complete adapted ID policy.

### v3

@@ Abstract
Heterogeneity poses a fundamental challenge for many real-world large-scale decision-making problems but remains largely understudied. In this paper, we study the fully heterogeneous setting of a prominent class of such problems, known as weakly-coupled Markov decision processes (WCMDPs). Each WCMDP consists of $N$ arms (or subproblems), which have distinct model parameters in the fully heterogeneous setting, leading to the curse of dimensionality when $N$ is large. We show that, under mild assumptions, an efficiently computable policy achieves an $O(1/\sqrt{N})$ optimality gap in the long-run average reward per arm for fully heterogeneous WCMDPs as $N$ becomes large. This is the first asymptotic optimality result for fully heterogeneous average-reward WCMDPs. Our main technical innovation is the construction of projection-based Lyapunov functions that certify the convergence of rewards and costs to an optimal region, even under full heterogeneity.
@@ 1 Introduction
Heterogeneity poses a fundamental challenge for many real-world decision-making problems, where each problem consists of a large number of interacting components. However, despite its practical significance, heterogeneity remains largely understudied in the literature. In this paper, we study heterogeneous settings of a prominent class of such problems, known as weakly-coupled Markov decision processes (WCMDPs) [23]. A WCMDP consists of $N$ arms (or subproblems), where each arm itself is a Markov decision process (MDP). In a heterogeneous setting, the MDPs could be distinct. At each time step, the decision-maker selects an action for each arm, which affects the arm’s transition probabilities and reward, and then the arms make state transitions independently. However, these actions are subject to a set of global budget constraints, where each constraint limits one type of total cost across all arms at each time step. The objective is to find a policy that maximizes the long-run average reward over an infinite time horizon. We focus on the planning setting, where all the model parameters (reward function, cost functions, budget and transition probabilities) are known.
WCMDPs have been used to model a wide range of applications, including online advertising [7, 55], job scheduling [48], healthcare [5], surveillance [40], and machine maintenance [21]. A faithful modeling of these applications calls for heterogeneity. For instance, in [5], arms are beneficiaries of a health program and they could react to interventions differently; in [40], arms are targets of surveillance who have different locations and probabilities to be exposed; in [21], arms are machines that could require distinct repair schedules.
Although heterogeneity is crucial in the modeling of these applications, most existing work on average-reward WCMDPs (or their special cases) establish asymptotic optimality only for the homogeneous setting where all arms share the same set of model parameters [42, 39, 14, 15, 25, 26, 27, 47, 24, 22]. The only known exceptions are [39] and [46]. They consider the typed heterogeneous setting, where the $N$ arms are divided into a constant number of types as $N$ scales up, with each type having distinct model parameters. Their budget constraints are also of special forms. While heterogeneous WCMDPs have been studied under the finite-horizon total-reward and discounted-reward criteria, as we review at the end of this section and also in Appendix A, these results do not extend to the average-reward setting we consider.
@@
Our contribution.
@@
In this paper, we study fully heterogeneous WCMDPs. We propose a policy we call the ID policy with reassignment, which generalizes the ID policy in the literature [26]. In our policy, we first perform an ID reassignment algorithm to reorder the arms, which ensures proper arm prioritization during policy execution. We then run a variant of the ID policy adapted to handle heterogeneity, which consists of two phases. The first phase is a pre-processing phase, where we compute an optimal single-armed policy for each arm (denoted as ${{\bar{\pi}}^{*}}_{i}$ for the $i$-th arm) that prescribes the ideal action the arm would take at each state. The second phase is the real-time phase. At each time step, the policy iterates over the arms according to their reassigned IDs, and it lets as many arms as possible follow their respective ideal actions while satisfying the budget constraints. Unlike the original ID policy, which has only one optimal single-armed policy due to homogeneous arms, our policy computes $N$ optimal single-armed policies, one for each arm. Our proposed policy is efficiently computable, with computational complexity polynomial in $N$.
We prove that the proposed ID policy with reassignment achieves an $O(1/\sqrt{N})$ optimality gap under mild assumptions as the number of arms $N$ becomes large. Here, the optimality gap refers to the difference between the long-run average reward per arm under our policy and that under the optimal policy. This is the first result establishing asymptotic optimality for fully heterogeneous average-reward WCMDPs.
We remark that the original ID policy was designed for a special case of WCMDPs known as restless bandits, and it is for the homogeneous setting. While the generalization in our proposed policy is natural, identifying the appropriate generalization and establishing its optimality gap in the fully heterogeneous setting are technically challenging and require new theoretical approaches.
Technical novelty.
@@
The main technical innovation of the paper is the introduction of a novel Lyapunov function for fully heterogeneous WCMDPs. Specifically, to prove the asymptotic optimality of a policy, a key step is to show that the system state is globally attracted to an optimal region where most arms can follow the ideal actions generated by their respective optimal single-armed policies ${{\bar{\pi}}^{*}}_{i}$’s. In the homogeneous setting, we can prove such convergence using state aggregation techniques that rely on the symmetry of arms. However, in the heterogeneous setting, states of different arms cannot be aggregated since arms are no longer symmetric. Our technique is to project arm states onto a set of carefully selected feature vectors, and define the Lyapunov function based on these projections. These feature vectors encode the minimal amount of information needed to evaluate the relevant functions of the system state (e.g., instantaneous reward or cost) and predict their future expectations. This projection-based Lyapunov function provides a principled way to measure deviations of the system state from the optimal region in a fully heterogeneous setting. A more detailed discussion of this approach can be found in Section 4.
@@
Beyond WCMDPs, our techniques have the potential to be applied to more general heterogeneous large stochastic systems. Heterogeneity has been a topic of strong interest in these systems, but it is known to be a challenging problem with limited theoretical results. Only recently have there been notable breakthroughs. [1, 2] extended the popular mean-field analysis to a class of heterogeneous large stochastic systems for the first time, but the results are only for transient distributions. Another line of work [54, 53] studied heterogeneous load-balancing systems. They first analyzed the transient distributions and then used interchange-of-limits arguments to extend the results to steady state. Our method provides a more direct framework for steady-state analysis and has the potential to generalize to a broader range of heterogeneous stochastic systems.
@@
We first focus on the average-reward criterion. As mentioned earlier, most existing work considers the homogeneous setting. Early work on WCMDPs primarily focuses on a special case known as the restless bandit (RB) problem, where each arm’s MDP has a binary action space (active and passive actions) and there is only one budget constraint that limits the total number of active actions across all arms at each time step. The seminal work by Whittle [44] introduced the RB problem and the celebrated Whittle index policy, which was later shown to achieve an $o(1)$ optimality gap as $N\to\infty$ under a set of conditions [42]. Subsequent work on RBs has focused on designing policies that achieve asymptotic optimality under more relaxed conditions [39, 25, 26, 47], as well as improving the optimality gap to $O(1/\sqrt{N})$ [25, 26] or $O(\exp(-cN))$ [14, 15, 27]. Among these papers, only [39] addresses heterogeneous RBs, but in the typed heterogeneous setting, where the $N$ arms are divided into a constant number of types as $N\to\infty$.
Beyond RBs, work on general average-reward WCMDPs is scarce. The closest to ours are [24, 39, 46], which studied WCMDPs with a single budget constraint and established $o(1)$ optimality gaps, but again in the homogeneous setting [24] or the typed heterogeneous setting [39, 46]. More recently, [22] proved the first $o(1)$ optimality gap result for WCMDPs with general budget constraints, but in the homogeneous setting.
Under the finite-horizon total-reward or discounted-reward criteria, there has been more work on heterogeneous settings, including both the typed heterogeneous setting [12, 17] and, more recently, the fully heterogeneous setting [8, 9, 10, 50]. However, the optimality gap in these papers generally grows super-linearly with the (effective) time horizon, except under restrictive conditions. As a result, it would be difficult to extend these results to the average-reward setting and still achieve asymptotic optimality.
Although our work focuses on the planning setting where all model parameters are known, there has been growing interest in developing reinforcement learning algorithms for the learning setting with unknown parameters [6, 33, 29, 32, 46, 30, 13, 36, 45, 3]. Many of these approaches rely on well-designed planning policies as a foundation to achieve learning efficiency. In this context, our results can serve as an important building block for developing model-based learning algorithms for fully heterogeneous WCMDPs.
@@
Let $\mathbb{R}$, $\mathbb{N}$, and $\mathbb{N}_{+}$ denote the sets of real numbers, nonnegative integers, and positive integers, respectively. Let $[N]\triangleq\{1,2,\dots,N\}$ for any $N\in\mathbb{N}_{+}$ and $[n_{1}:n_{2}]\triangleq\{n_{1},n_{1}+1,\dots,n_{2}\}$ for $n_{1},n_{2}\in\mathbb{N}_{+}$ with $n_{1}\leq n_{2}$. Let $[0,1]_{N}=\{i/N\colon i\in\mathbb{N},0\leq i/N\leq 1\}$, the set of integer multiples of $1/N$ in $[0,1]$. For a matrix ${A}\in\mathbb{R}^{d\times d}$, we denote its operator norm as $\left\lVert{A}\right\rVert_{p}=\sup_{x\neq\mathbf{0}}\frac{\left\lVert{A}x\right\rVert_{p}}{\left\lVert x\right\rVert_{p}}$. We use boldface letters to denote matrices, and regular letters to denote vectors and scalars. We write $\mathbb{R}^{\mathbb{S}}$ for the set of real-valued vectors indexed by elements of $\mathbb{S}$, or equivalently, the set of real-valued functions on $\mathbb{S}$; for each $v\in\mathbb{R}^{\mathbb{S}}$, let $v(s)$ to denote its element corresponding to $s\in\mathbb{S}$.
@@ 2 Problem setup
A policy $\pi$ for the $N$-armed problem specifies the action for each of the $N$ arms, in a possibly history-dependent way. Under policy $\pi$, let $S_{i,t}^{\pi}$ denote the state of the $i$th arm at time $t$, and we refer to $\bm{S}^{\pi}_{t}\triangleq(S_{i,t}^{\pi})_{i\in[N]}$ as the system state. Similarly, let $A_{i,t}^{\pi}$ denote the action applied to arm $i$ at time $t$, and we refer to $\bm{A}^{\pi}_{t}\triangleq(A_{i,t}^{\pi})_{i\in[N]}$ as the system action. In this paper, we also use an alternative representation of the system state, denoted as ${\bm{X}}_{t}^{\pi}$ and defined as follows. Let $X_{i,t}^{\pi}=(X_{i,t}^{\pi}(s))_{s\in\mathbb{S}}\in\mathbb{R}^{|\mathbb{S}|}$ be a row vector where the entry corresponding to state $s$ is given by $X_{i,t}^{\pi}(s)=\mathbbm{1}\{S_{i,t}^{\pi}=s\}$; i.e., $X_{i,t}^{\pi}$ is a one-hot row vector whose $s$’s entry is $1$ if $S_{i,t}^{\pi}=s$ and is 00 otherwise. Then let ${\bm{X}}_{t}^{\pi}$ be an $N\times|\mathbb{S}|$ matrix whose $i$th row is $X_{i,t}^{\pi}$. It is easy to see that ${\bm{X}}_{t}^{\pi}$ contains the same information as $\bm{S}^{\pi}_{t}$, and we refer to both of them as the system state. In this paper, we often encounter vectors like $X_{i,t}^{\pi}=(X_{i,t}^{\pi}(s))_{s\in\mathbb{S}}$, whose entries correspond to different states in $\mathbb{S}$. For such vectors, say $u$ and $v$, we use the inner product to write a sum for convenience $\left\langle u,v\right\rangle\triangleq\sum_{s\in\mathbb{S}}u(s)v(s)$. We sometimes omit the superscript $\pi$ when it is clear from context.
@@
Our objective is to maximize the long-run time-average reward subject to the cost constraints. To be more precise, we follow the treatment for maximizing average reward in [35]. For any policy $\pi$ and an initial state $\bm{S}_{0}$ of the $N$-armed system, consider the limsup average reward $R^{+}(\pi,\bm{S}_{0})$ and the liminf average $R^{-}(\pi,\bm{S}_{0})$, defined as $R^{+}(\pi,\bm{S}_{0})=\lim\sup_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{N}\sum_{i\in[N]}\mathbb{E}\left[r_{i}(S_{i,t}^{\pi},A_{i,t}^{\pi})\right]$ and $R^{-}(\pi,\bm{S}_{0})=\lim\inf_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\frac{1}{N}\sum_{i\in[N]}\mathbb{E}\left[r_{i}(S_{i,t}^{\pi},A_{i,t}^{\pi})\right]$. If $R^{+}(\pi,\bm{S}_{0})=R^{-}(\pi,\bm{S}_{0})$, then the average reward of policy $\pi$ under initial condition $\bm{S}_{0}$ exists and is defined as
@@
Let the optimal value of this problem be denoted as $R^{*}(N,\bm{S}_{0})$. Note that since the WCMDP is an MDP with finite state and action space, if we replace the $R^{-}(\pi,\bm{S}_{0})$ in the objective (3a) with $R^{+}(\pi,\bm{S}_{0})$, the optimal value stays the same [35, Proposition 9.1.6].
@@
Recall that exactly solving the WCMDP problem is PSPACE-hard [34]. In this paper, our goal is to design a policy $\pi$ that is efficiently computable and asymptotically optimal as $N\to\infty$, with the following notion for asymptotic optimality. For any policy $\pi$, we define its optimality gap as $R^{*}(N,\bm{S}_{0})-R^{-}(\pi,\bm{S}_{0})$. We say the policy $\pi$ is asymptotically optimal if as $N\to\infty$,
@@
When we take this asymptotic regime as $N\to\infty$, we keep the number of constraints, $K$, as well as the budget coefficients, $\alpha_{1},\alpha_{2},\dots,\alpha_{K}$, fixed. We assume that the reward functions and cost functions are uniformly bounded, i.e., $\sup_{i\in\mathbb{N}_{+}}\max_{s\in\mathbb{S},a\in\mathbb{A}}|r_{i}(s,a)|\triangleq r_{\max}<\infty$ and $\sup_{i\in\mathbb{N}_{+}}\max_{k\in[K],s\in\mathbb{S},a\in\mathbb{A}}c_{k,i}(s,a)\triangleq c_{\max}<\infty$. This notion for asymptotic optimality is consistent with that in the existing literature (e.g., [39, Definition 4.11]). We are interested in not only achieving asymptotic optimality but also characterizing the order of the optimality gap.
@@
$\sum_{s^{\prime}\in\mathbb{S},a^{\prime}\in\mathbb{A}}\mathbb{P}_{i}(s\mid s^{\prime},a^{\prime})y_{i}(s^{\prime},a^{\prime})=\sum_{a\in\mathbb{A}}y_{i}(s,a),\;\forall s\in\mathbb{S},\forall i\in[N],$ (5c)
$\sum_{s^{\prime}\in\mathbb{S},a^{\prime}\in\mathbb{A}}y_{i}(s^{\prime},a^{\prime})=1,\;y_{i}(s,a)\geq 0,\;\forall s\in\mathbb{S},\forall a\in\mathbb{A},\forall i\in[N].$ (5d)
@@
Therefore, to derive an upper bound for the optimality gap, it suffices to control $R_{N}^{\textup{rel}}-R^{-}(\pi,\bm{S}_{0})$, which we will show is $O(1/\sqrt{N})$ in Theorem 1.
@@
Our LP (5) serves a similar role to the LP used in previous work on restless bandits and WCMDPs with homogeneous arms [see, e.g., 42, 14, 25]. Both our LP and the LP in previous work relax the hard budget constraints to time-average constraints. However, in the homogeneous arm setting, the LP has only one set of state-action frequencies $y(s,a)$, and the LP is independent of $N$. As a result, both the optimal value of the LP and the complexity of solving it are independent of $N$. Some existing work [39] considers heterogeneous arms, but only in the limited sense of having a constant number of arm types. This setting still closely resembles the homogeneous setting, and the LP remains independent of $N$.
@@
3 ID policy with reassignment
In this section, we introduce the ID policy with reassignment, generalized from the ID policy designed for homogeneous restless bandits in the literature [26]. Our policy first performs an ID reassignment procedure, and then proceeds to run a variant of the ID policy adapted to handle heterogeneity. We begin by introducing a building block of our policy, referred to as optimal single-armed policies, followed by the ID reassignment algorithm and the execution of the adapted ID policy.